an electicial contractor paid $1040 for 36 light fixtures to be installed in a hotel lobby. if the first type of fixture costs $25 each and the second type of fixture costs $36 each, how many of each kind did he buy?
Answer by jorel1380(3719)

n+p=36
25n+36p=1040
25n+25p=900
11p=140
p=12.72
n=23.28
The contractor bought 23.28 of the $25 fixture, and 12.72 of the $36 fixture..